OpenStudy (anonymous):

Find the surface area and volume of a sphere with a radius of 8cm. Use pi equals 3.14 and round to the nearest hundredth. can someone help me please

OpenStudy (anonymous):

the equation for surface area of a sphere is SA = 4(pi)r^2 if we know that the radius (r) is 8cm, we can plug values in for the Surface Area equation :) SA = 4(3.14)(8)^2 = 803.84 The surface area of the sphere is 803.84 cm^2 hope that helped! :)

OpenStudy (anonymous):

can you help find the volume

OpenStudy (anonymous):

yep!! volume = (4/3)(pi)(r)^3 plugging in these values again, V = (4/3)(3.14)(8)^3 = 2,143.57 The volume is 2143.57 cm^3

OpenStudy (goldphenoix):

To find the volume of a sphere, you must use the formula: \[\large \large V = \frac{ 4 }{ 3 } × \pi × r^3\] We know that the radius is 8. So cube it to get 512. Next, you times it by 4/3 to get 682.67. Final step is to pi it. The volume should be 2144.66 m^3.

OpenStudy (anonymous):

thank you